A used car is on sale for a discount of 20% off the regular price of P550 000. an additional 10% discount on the sale price was offered. Is the result a 30% discount? What is the single discount that would give the same sale price?
please give me your representation thank you!!!!

Answer by Mathtut(3670)

no, the result is not a 30% discount...why because you are taking 20% on the original price and an addition 10% on the sales price.....if you were taking both discounts on the original price then the answer would be yes.
550000(.2)=110000 so the sales price is 440000(550000-110000) for which we take another 10% off 440000(.1)=44000. so the final sales price is 440000-44000=396000.
the final discount percentage can be found by taking the total of discounts 110000+44000=154000 and divide it by the orginal cost 550000
:
154000/550000=.28 or 28% discount